June 18, 2013 Calorimetry
'Teleconference Time: 11:00 a.m. EST
- ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
- Phone connection only upon request.
- +1-866-740-1260 : US+Canada
- +1-303-248-0285 : International
- then enter participant code: 3421244# (remember the "#").
- or www.readytalk.com (and code without the #)
Participant Direct Lines
- JLab Phone: in CC F326 is 757-269-6460 (usual room)
- JLab Phone in CC L207 is 757-269-7084
- Phone in the Regina Videoconference Suite is 306-585-4204
- Athens Phone: in Christina's office is 011-30-210-727-6947
Items for followup from previous meeting(s)
- Light guides
- Shipment of replacement light guides (USM)
- Action Items
- Light guide gluing
- SiPM studies
- Cosmics analysis
- Decommissioning of Bcal module factory
- Reconstruction and Simulation
- TDR update
- Calibration of Calorimeters
- Any other business
Attendees: David, Beni, John L., Serguei (JLab); Alam Toro, Elias Rojas, Will Brooks (Chile); GV2, Pavlos, Stratos (Athens); Shaun, Zisis (Regina)
- Announcements: none
- Action Items: Chile responded offline to Elton's request on an update regarding the light guide re-machining.
- Monitoring: Stratos responded offline to Elton's question regarding the FCAL plexiglas sheet specs. G. Vassileiadis will meet with David and mark at JLab on software projects.
- FCAL (John L.): The dark room is at JLab. Rectangular structure is in place. Some metal skin has been mounted but mostly awaiting to put cables in first; 4x4 ft^2 skin pieces go on in an hour, using rivets. Working on dark room sensors (humid., temp., pressure, light) and debugging them. Labyrinth delivered to IU, awaiting some sheet metal, and will be shipped to JLab late this week or next. On the JLab website, the FCAL is (finally) shown as completed! (Yippie).
- Light guide gluing (from Elton):
- Rework of all light guides was completed last week, including illumination of guides that appeared to be fine to ensure complete curing
- Preparations for gluing of last 13 modules is completed. This included
- sanding light guides
- obtaining a complete count of non-conforming light guides (35 additional to the 109 on those returned to USM). I will send out a complete log later today.
- preparing cooling plates for inclusion into the readout assembly
- covers for remaining modules
- Gluing of the remaining 13 new modules will begin today. Laura and Tegan will be trained so they can help out with the gluing.
- SiPM studies: IV curve measurements as a function of temperature continue at Regina. They should be completely soon.
- Cosmics analysis. Shaun presented his cosmics analysis (new stuff starts on slide 19) and discussion ensued. The scripts are ready for use on all module cosmics files and a few issues (some small systematic shifts in the U/D ratio for modules 29 and 39 (versus the golden runs of module 26) and the 1:4 ratio. Beni suggested that Shaun plots (ratio_mod - ratio_golden)/ratio_golden to show differences in percent.
- Decommissioning of Bcal module factory: this will proceed in August. If something happens that delays the BCAL installation schedule, we could change this but that would require an extension to the contract as it expires on August 31.
- Light guide gluing (from Elton):
- Reconstruction and Simulation: nothing to report.
- TDR update: nothing to report.
- Calibration of Calorimeters: nothing to report.
- Any other business: none.